H_{3}^{+} ABUNDANCES MODELED WITH A VARIABLE COSMIC RAY IONIZATION RATE

\underline{PAUL B. RIMMER}, ERIC HERBST, Physics Department, The Ohio State University, Columbus, OH 43202.

Recent observations of H_{3}^{+} abundances in diffuse clouds have suggested a higher value for the cosmic ray ionization than has been generally accepted. We provide a model that takes into account a variable cosmic ray ionization, and show that it produces H_{3}^{+} abundances close to the observed values. We also apply this method to an IR-dense cloud model.